lbw  1230-1470 filter (filter number 3)


        The lbw cal measurements have shown a jump in calDefl/Tsys for polB when the 1230 to 1470 Mhz filter is installed. To test this the hcorcal was fired using the 1 second toggle for about 2400 seconds with the 1230-1470 filter in (1 second cal on , 1 sec cal off). This was followed by 1200 seconds of calon,off with the 1230-1470 Mhz filter out. The telescope was sitting at az=280, za=10 for these measurements. The lst was around a hours.

Data reduction:

    The data reduction consisted of:

The data

    The plots show Tsys versus time (.ps)  (.pdf):     A Tsys measurement was made every 2 seconds (1 sec calon, 1 sec caloff).  The positive going spikes are sources drifting through the beam. The blue 1425 Mhz Tsys is about 7% less then than the 1350,1375, and 1400 Mhz Tsys.  This may be a problem in the calvalues in use (if so, the error in in both polA and polB). The transition from filter to no filter at 14.14 hours is smooth for polA. For polB the 1425 Mhz (blue) tsys jumps down. This is not the 1230-1470 filter since it would affect all 4 of the bands.

If the problem is the filter bank relay not closing completely a better test would be to take 20 seconds of filter in, 20 seconds of filter off , and then repeat this for a few tens of cycles.